a wooden cube just floats inside water when a 200g mass is placed on it. when the mass is just removed, the cube is 2cm above the water level. the size of cube is

Grade:10

3 Answers

sakshi gupta
8 years ago
Let length of one edge = x
Vol. of cube = L*B*W
Vol. of cube = x*x*x
Vol. of cube = x^3
Vol. of cube = Area * height.
Vol. of cube = x² * 2
Vol. of cube = 2x² cm^3
200 mg = 2x²
x² = 200 mg / 2
x = | ± √100 |
x = 10 cm
Alia Bhatt
6 years ago
Hello Ian Alia Bhatt@shanaya@humpty Sharma ki dulhania@highway@dear zindagi girl@kapur and sons ki Dulhan@badrinath ki Dulhan isCall 9414738393
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Let length of one edge = x Vol. of cube
= L*B*W Vol. of cube
= x*x*x Vol. of cube
= x^3
Vol. of cube = Area * height.
Vol. of cube = x² * 2
Vol. of cube = 2x² cm^3
200 mg = 2x²
x² = 200 mg / 2
x = | ± √100 |
x = 10 cm

Thanks and Regards